OpenStudy (anonymous):

write the difference as a mixed #. simplify answer. -2 1/6 - (-4 1/2)

OpenStudy (anonymous):

\[-2\tfrac{1}{6}=-\frac{13}{6}\] \[-(-4\tfrac{1}{2})=\frac{9}{2}\] so your job is \[\frac{9}{2}-\frac{13}{6}\]

OpenStudy (anonymous):

you can write the first fraction as an equivalent fraction with denominator 6 by multiplying top and bottom by 3 to get \[\frac{27}{6}-\frac{13}{6}\] and now it is easy

OpenStudy (anonymous):

so now you just have to subtract 17 & 13?

OpenStudy (anonymous):

i get \[\frac{14}{6}=\frac{7}{3}=2\frac{1}{3}\]

jhonyy9 (jhonyy9):

-13/6 - (-9/2) = 9/2 - 13/6

OpenStudy (anonymous):

i mean 27-13

OpenStudy (anonymous):

@jaycyy subtract 13 from 27 i.e. 27-13

OpenStudy (anonymous):

yes that is it. keep the denominator at 6

jhonyy9 (jhonyy9):

yes (27-13)/6

OpenStudy (anonymous):

then because you were told you have to turn the answer back in to a mixed number.

OpenStudy (anonymous):

is it 13/6

OpenStudy (anonymous):

no try 27-13 again

OpenStudy (anonymous):

i ment 14/6

OpenStudy (anonymous):

yes that is the one

OpenStudy (anonymous):

do yu simplify

jhonyy9 (jhonyy9):

what equal 7/3

OpenStudy (anonymous):

in fact you have to do more than reduce. you have to turn it back into a mixed number

OpenStudy (anonymous):

i dont think i knoww :(

OpenStudy (anonymous):

i bet you actually do lets go slow

OpenStudy (anonymous):

okkk

OpenStudy (anonymous):

we have \[\frac{14}{6}\] first of all. now can you reduce this fraction?

OpenStudy (anonymous):

by which i mean reduce to lowest terms, not write as a mixed number. we will do that second

OpenStudy (anonymous):

7/3

OpenStudy (anonymous):

ok one step away

OpenStudy (anonymous):

how many times does 3 go in to 7?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

remainder is ?

OpenStudy (anonymous):

clear what i am asking? 3 goes in to 7 2 times. the remainder is 1, so mixed number is \[2\tfrac{1}{3}\]